Assume that the kinetic energy of a 1350 kg car moving at 135 km/hr could be converted entirely into heat. What amount of water could be heated from 23 (degrees C) to 45 (degrees C) by the car's energy?
Answer #1
example

Kinetic Energy (KE)= 1/2 mv^2
where m=mass in kg, v=velocity in meters/sec and KE is in Joules
convert 145 km/hr to m/s=40.3 meters/sec (see link to conversion calculator, or you can convert manually)
KE= 0.5 * 1050kg * (40.3)^2=852.6kJ
The heat capacity of water (Cp) at 25C is 4.2 J/gm-K (wikipedia)
Use the formula Q=mCpdT to calculate the mass in grams
Q= heat in Joules, m=mass in grams, dT= delta T in K
dT= 36K
m=Q/Cp/dT
m=852600 J/4.2/36=5638.9 gms